# GLEE: A Unified Framework and Benchmark for Language-based Economic Environments

### v1

@@ Appendix E Graphs
$\delta_{A}$ Figure 14 $F_{A}$ Figure 20 p Figure 26
$\delta_{B}$ Figure 15 $F_{B}$ Figure 21 v Figure 27
$M$ Figure 16 $M$ Figure 22 $M$ Figure 28
$T$ Figure 17 $T$ Figure 23 Buyer type Figure 29
CI Figure 18 CI Figure 24 CI Figure 30
MA Figure 19 MA Figure 25 Messages type Figure 31
Table 6: The table summarizes the various figures in the appendix and serves as a table of contents for easy access to them. Each figure presents the impact of changing parameter values on the different metrics. CI = Complete Information. MA = Textual messages allowed.
E.1 Bargaining
Figure 14: The effect of the change in parameter $\delta_{A}$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in bargaining games. The change is measured relative to the scenario where $\delta_{A}$ is $0.8$ (i.e. $\Delta$ = (Metric |||| $\delta_{A}$ = $NewValue$) - (Metric |||| $\delta_{A}$ = $0.8$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 15: The effect of the change in parameter $\delta_{B}$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in bargaining games. The change is measured relative to the scenario where $\delta_{B}$ is $0.8$ (i.e. $\Delta$ = (Metric |||| $\delta_{B}$ = $NewValue$) - (Metric |||| $\delta_{B}$ = $0.8$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 16: The effect of the change in parameter $M$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in bargaining games. The change is measured relative to the scenario where $M$ is $100$ (i.e. $\Delta$ = (Metric |||| $M$ = $NewValue$) - (Metric |||| $M$ = $100$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 17: The effect of the change in parameter $T$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in bargaining games. The change is measured relative to the scenario where $T$ is $12$ (i.e. $\Delta$ = (Metric |||| $T$ = $NewValue$) - (Metric |||| $T$ = $12$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 18: The effect of the change in parameter Complete Information on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in bargaining games. The change is measured relative to the scenario where Complete Information is $False$ (i.e. $\Delta$ = (Metric |||| Complete Information = $NewValue$) - (Metric |||| Complete Information = $False$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 19: The effect of the change in parameter Messages Allowed on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in bargaining games. The change is measured relative to the scenario where Messages Allowed is $False$ (i.e. $\Delta$ = (Metric |||| Messages Allowed = $NewValue$) - (Metric |||| Messages Allowed = $False$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
E.2 Negotiation
Figure 20: The effect of the change in parameter $F_{A}$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in negotiation games. The change is measured relative to the scenario where $F_{A}$ is $0.8$ (i.e. $\Delta$ = (Metric |||| $F_{A}$ = $NewValue$) - (Metric |||| $F_{A}$ = $0.8$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 21: The effect of the change in parameter $F_{B}$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in negotiation games. The change is measured relative to the scenario where $F_{B}$ is $0.8$ (i.e. $\Delta$ = (Metric |||| $F_{B}$ = $NewValue$) - (Metric |||| $F_{B}$ = $0.8$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 22: The effect of the change in parameter $M$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in negotiation games. The change is measured relative to the scenario where $M$ is $100.0$ (i.e. $\Delta$ = (Metric |||| $M$ = $NewValue$) - (Metric |||| $M$ = $100.0$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 23: The effect of the change in parameter $T$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in negotiation games. The change is measured relative to the scenario where $T$ is $1$ (i.e. $\Delta$ = (Metric |||| $T$ = $NewValue$) - (Metric |||| $T$ = $1$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 24: The effect of the change in parameter Complete Information on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in negotiation games. The change is measured relative to the scenario where Complete Information is $False$ (i.e. $\Delta$ = (Metric |||| Complete Information = $NewValue$) - (Metric |||| Complete Information = $False$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 25: The effect of the change in parameter Messages Allowed on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in negotiation games. The change is measured relative to the scenario where Messages Allowed is $False$ (i.e. $\Delta$ = (Metric |||| Messages Allowed = $NewValue$) - (Metric |||| Messages Allowed = $False$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
E.3 Persuasion
Figure 26: The effect of the change in parameter $p$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in persuasion games. The change is measured relative to the scenario where $p$ is $0.333333333$ (i.e. $\Delta$ = (Metric |||| $p$ = $NewValue$) - (Metric |||| $p$ = $0.333333333$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 27: The effect of the change in parameter $v$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in persuasion games. The change is measured relative to the scenario where $v$ is $1.2$ (i.e. $\Delta$ = (Metric |||| $v$ = $NewValue$) - (Metric |||| $v$ = $1.2$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 28: The effect of the change in parameter $M$ on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in persuasion games. The change is measured relative to the scenario where $M$ is $100$ (i.e. $\Delta$ = (Metric |||| $M$ = $NewValue$) - (Metric |||| $M$ = $100$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 29: The effect of the change in parameter Buyer is Myopic on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in persuasion games. The change is measured relative to the scenario where Buyer is Myopic is $False$ (i.e. $\Delta$ = (Metric |||| Buyer is Myopic = $NewValue$) - (Metric |||| Buyer is Myopic = $False$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 30: The effect of the change in parameter Complete Information on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in persuasion games. The change is measured relative to the scenario where Complete Information is $False$ (i.e. $\Delta$ = (Metric |||| Complete Information = $NewValue$) - (Metric |||| Complete Information = $False$)), and is presented for each of the models that played Alice (left column) and Bob (right column).
Figure 31: The effect of the change in parameter Message Type on the metrics Self Gain (top row), Efficiency (middle row), and Fairness (bottom row), in persuasion games. The change is measured relative to the scenario where Message Type is $binary$ (i.e. $\Delta$ = (Metric |||| Message Type = $NewValue$) - (Metric |||| Message Type = $binary$)), and is presented for each of the models that played Alice (left column) and Bob (right column).

### v2

@@ F.2 Baseline Models
$\delta_{A}$ 0.9 $F_{A}$ 1 p 0.5
$\delta_{B}$ 0.9 $F_{B}$ 1 v 1.25
$M$ $10^{4}$ $M$ $10^{4}$ M $10^{4}$
$T$infty $T$ 1 $T$ 20
CI True CI True CI True
MA False MA False Messages type Binary
Buyer type Long-living
F.3 Predictive Validity of Linear Regression
Section 4 describes the method used for our analyses, based on beta coefficient interpretation in linear regression. In this appendix, we demonstrate that linear regression achieves strong predictive performance compared to state-of-the-art (SOTA) models on tabular regression tasks.
For each of the three game families (bargaining, negotiation, persuasion) and each of the four evaluation metrics (Efficiency, Fairness, Alice’s self-gain, Bob’s self-gain), we trained XGBoost ([11]) and CatBoost ([35]) as SOTA baselines.
The dataset was randomly split into 80% training and 20% testing. All models received the full feature set as input (see §4 for details) and were tasked with predicting the target metric.
Table 9 summarizes the results. We find that the linear model performs on par with, and occasionally outperforms, the more complex SOTA models, with only minor differences in RMSE. These findings suggest that the relationships between game parameters and outcomes are largely linear, and thus support the use of beta coefficients as a reliable and interpretable tool for analyzing how input features, such as game configurations and agent identities, influence performance metrics.
Table 9: Root Mean Squared Error (RMSE) for each model across families and metrics. Values are reported as mean ± standard deviation over 100 random seeds.
Family Metric Linear Regression XGBoost CatBoost
Bargaining Alice Self Gain 0.134 ± 0.001 0.134 ± 0.001 0.134 ± 0.001
Bob Self Gain 0.129 ± 0.002 0.129 ± 0.002 0.129 ± 0.002
Efficiency 0.131 ± 0.002 0.131 ± 0.002 0.131 ± 0.002
Fairness 0.157 ± 0.003 0.156 ± 0.003 0.156 ± 0.003
Negotiation Alice Self Gain 0.137 ± 0.014 0.137 ± 0.014 0.136 ± 0.014
Bob Self Gain 0.139 ± 0.013 0.138 ± 0.012 0.138 ± 0.012
Efficiency 0.333 ± 0.004 0.331 ± 0.004 0.331 ± 0.004
Fairness 0.090 ± 0.003 0.088 ± 0.003 0.088 ± 0.003
Persuasion Alice Self Gain 0.356 ± 0.003 0.353 ± 0.003 0.352 ± 0.003
Bob Self Gain 0.834 ± 0.011 0.842 ± 0.011 0.840 ± 0.011
Efficiency 0.388 ± 0.004 0.380 ± 0.004 0.378 ± 0.004
Fairness 0.379 ± 0.004 0.372 ± 0.005 0.371 ± 0.004
Appendix G Additional Results
G.1 LLMs Performance (Q2)
In this appendix, we present additional result graphs that complement those shown in §4.
In this sub-appendix, we present graphs that help address Q2: How do different LLMs behave in strategic interactions, and which models achieve fair, efficient, and high self-gain outcomes?
Figures 21, 22, and 23 show how the pair of models playing bargaining, negotiation, and persuasion games influenced the efficiency and fairness of the game.
(a) Efficiency
1 2 3 4 5 6 7 8 9 A B C D
1. claude-3-5-sonnet
2. claude-3-7-sonnet
3. gemini-1.5-flash
4. gemini-1.5-pro
5. gemini-2.0-flash
6. gemini-2.0-flash-lite
7. gpt-4o
8. gpt-4o-mini
9. grok-2-1212
A. llama-3.1-405b
B. llama-3.3-70b
C. mistral-large-2411
D. o3-mini
0.210.21-0.21- 0.210.050.05-0.05- 0.050.110.110.110.11
(b) Fairness
1 2 3 4 5 6 7 8 9 A B C D
1.
2.
3.
4.
5.
6.
7.
8.
9.
A.
B.
C.
D.
0.240.24-0.24- 0.240.10.1-0.1- 0.10.050.050.050.05
Figure 21: The effect of the identities of the two players (rows: Alice, columns: Bob) on efficiency and fairness in Bargaining games, reported relative to the mean outcome.
(a) Efficiency
1 2 3 4 5 6 7 8 9 A B C D
1. claude-3-5-sonnet
2. claude-3-7-sonnet
3. gemini-1.5-flash
4. gemini-1.5-pro
5. gemini-2.0-flash
6. gemini-2.0-flash-lite
7. gpt-4o
8. gpt-4o-mini
9. grok-2-1212
A. llama-3.1-405b
B. llama-3.3-70b-
C. mistral-large-2411
D. o3-mini
0.130.13-0.13- 0.130.010.01-0.01- 0.010.10.10.10.1
(b) Fairness
1 2 3 4 5 6 7 8 9 A B C D
1.
2.
3.
4.
5.
6.
7.
8.
9.
A.
B.
C.
D.
0.070.07-0.07- 0.070.030.03-0.03- 0.030.020.020.020.02
Figure 22: The effect of the identities of the two players (rows: Alice, columns: Bob) on efficiency and fairness in Negotiation games, reported relative to the mean outcome.
(a) Efficiency
1 2 3 4 5 6 7 8 9 A B C D
1. claude-3-5-sonnet
2. claude-3-7-sonnet
3. gemini-1.5-flash
4. gemini-1.5-pro
5. gemini-2.0-flash
6. gemini-2.0-flash-lite
7. gpt-4o
8. gpt-4o-mini
9. grok-2-1212
A. llama-3.1-405b
B. llama-3.3-70b
C. mistral-large-2411
D. o3-mini
0.330.33-0.33- 0.330.040.04-0.04- 0.040.240.240.240.24
(b) Fairness
1 2 3 4 5 6 7 8 9 A B C D
1.
2.
3.
4.
5.
6.
7.
8.
9.
A.
B.
C.
D.
0.440.44-0.44- 0.440.080.08-0.08- 0.080.280.280.280.28
Figure 23: The effect of the identities of the two players (rows: Alice, columns: Bob) on efficiency and fairness in Persuasion games, reported relative to the mean outcome.
Appendix H Ethics Statement
Furthermore, our framework demonstrates the capability of collecting data from human players to understand the differences and similarities between LLMs and humans in economic environments. While this line of research has the potential for a strong scientific contribution, particularly in the field of behavioral economics, it also raises several ethical considerations. The collection of human data must be conducted with careful regulation and adherence to clear ethical guidelines. The entire process of data collection from human participants is elaborated in Appendix E.